# General-Purpose Multicore Architectures

### v1

@@ 1 Introduction
In this chapter, we will examine six topics related to multicore CPU architectures. First, we will motivate the benefits and limitations of processing multiple tasks concurrently, and how they drove the need for parallel processing. Second, we will examine the hardware design of a typical multicore CPU, and the changes required to support the efficient execution of multiple programs on multiple CPU cores. Third, we will study how memory management policies change to handle the increased traffic from multiple cores. Fourth, we will address software issues that optimize the use of multicore CPUs. Fifth, we will introduce common metrics that are used in the context of multicore CPUs. Finally, we will close by briefly discussing how commercial multicore CPUs have evolved since their introduction.
@@ 2 Motivating the Need for Concurrent Processing
As early as the days of analog computing, the benefits of parallelizing tasks became a clear goal: Luigi Federico Menabrea, in his 1842 study of Charles Babbage’s Analytical Engine [59], noted that a key advantage of mechanized computation would be its ability to produce several results at the same time. In the mid 20th century, as digital computers evolved rapidly, there was a pressing need to maximize both the performance and utilization of these computers, and the extraction of parallelism, in its various forms, became a key approach to meet these needs. While we focus on multicore CPUs in this chapter, we provide a brief taxonomy of parallel computing hardware in Section 2.1.
@@ 2.1 Classifying Parallel Computing Hardware
Under Flynn’s taxonomy, non-pipelined, non-superscalar single-core CPUs that employ a von Neumann architecture are a canonical example of a SISD architecture. Examples of SIMD architectures include graphics processing units (GPUs), array processors, and vector extensions for CPUs. One reasonably-accepted example of a MISD architecture is redundant computing, such as the systems employed in avionics units. [^3] There is some latent controversy about which category pipelined CPUs fall into. In our best imitation of Switzerland and its neutrality, we do not take a stand for any category. Sorry for being so bland. The MIMD category includes most computers that are capable of multiprocessing (Section 2.2), including computers with multicore CPUs.
In a multicore CPU, each CPU core is capable of executing one or more independent streams of instructions, with each instruction stream operating on its own data set (though this data set can potentially overlap with the data sets being used by other cores). To support this MIMD model of execution, each core has its own control logic, and much of the computation takes place independent from other cores. We elaborate on this more in Section 3, while also discussing when coordination between cores does occur (both explicitly and implicitly).
@@ 2.3 Thread-Level Parallelism (TLP) Within an Application
where $f$ is the fraction of the program that can be parallelized, and $n$ is the number of concurrent threads that can execute the parallelizable part of the program. To visualize Amdahl’s Law, we show an example in Figure 3 where $f=0.4$ (i.e., 40% of the program can be parallelized), for various thread counts. As the figure shows, for every doubling of the thread count, the reduction in total time is only half of the reduction achieved by the previous doubling.
@@
Putting this all together, we show a synthetically-constructed example of the parallel speedup one could expect to observe on a real parallel system in Figure 5. In our example, we assume that 99% of the application can be parallelized across all threads, while the remaining 1% must execute sequentially. We also assume that the machine has 32 CPU cores, and since it has a fixed set of available resources, we use Amdahl’s Law to predict performance. We make three observations from the figure. First, even with a single thread, the observed speedup is less than 1. This is because parallel speedup is compared against the best sequential implementation (see Section 6), and a parallel implementation (with a configurable parameter $n$ that sets the thread count) typically has overheads associated with adding code to support parallel execution, and these overheads are observed when $n=1$. Second, super-linear speedups can even exceed perfect parallelism. As mentioned above, this is the result of threads helping each other through data locality. Third, after a certain thread count (24 in our example), the performance at a higher thread count is actually lower than the performance at a lower thread count. This is because the benefits of additional thread-level parallelism and additional data locality are overcome by synchronization and interference overheads.
@@ 2.4 What to Do With All These Transistors?
Through the 1990s (and drawing some very broad, and likely inaccurate, generalizations), the vast majority of parallel computing architectures catered to the large-scale computing market, with much of the effort focused on supercomputers. Most personal computers (PCs) and workstations [^6] At that time, there was a distinction between more mainstream PCs and high-performance workstations that targeted power users. Today, in large part due to advancements in PC capabilities throughout the 1990s, this distinction has mostly disappeared, but we reference it here for historical context. incorporated only a single-core CPU (or in the case of some high-end workstations, multiple single-core CPUs installed on a multi-socket motherboard). Much of the focus on architectural innovation remained on improving the performance of the single core, leading to the maturation and widespread commercial adoption of techniques such as out-of-order processing and superscalar execution.
@@
As the breakdown of Dennard scaling materialized during the first few years of the 2000s, the power consumption issue was exacerbated significantly. This added pressure led the next wave of multicore CPUs to make a key trade-off to avoid the impending limits of thermal density: they operated each CPU core at a lower frequency than their single-core contemporaries, in the hopes that the increased parallel processing capability would make up for reduced single-core performance [20]. We discuss more about the CPU core trade-offs in Section 3.1, which allowed architects to harness TLP instead of aggressive ILP to improve performance while keeping power issues under control, and opened the floodgates for the proliferation of multicore CPUs.
@@ 3.1 Optimizing CPU Cores for Parallelism
Multicore CPU designers prevent the design and verification of the CPU from increasing linearly with the core count by relying on design reuse through tiling. In a homogeneous multicore CPU, the design of the cores is identical, in order to reduce manufacturing and verification/testing costs. (We discuss popular approaches for heterogeneous multicore CPU design in Section 7.) Figure 6 shows how the majority of the chip for a typical homogeneous multicore CPU is made up of multiple tiles, where each tile has exactly the same components. A tile usually consists of a single CPU core and the private caches associated with that core. Some architectures include a slice of the last-level cache in the tile (see Section 3.2). In addition to these tiles, the chip includes shared components (e.g., controllers that connect the CPU cores to external components, interconnects).
@@ 3.2 Sharing Caches and Main Memory
One issue with sharing resources across CPU cores is interference [6], where due to the finite amount of available resources, multiple cores may contend with each other for their desired share of these resources. Let us look at a contrived example, with a two-core CPU, where one core is servicing a cache miss that must retrieve data from main memory. As the data for that miss arrives at the LLC, the LLC must evict an existing cache block (assuming that no invalid blocks remain). Interference can occur when the eviction is for a cache block belonging to the other core, as that core will now miss on a subsequent access to the evicted block. Had the first core not triggered the eviction, this subsequent access would have been a cache hit, avoiding the long-latency miss to main memory. The resource contention that results from multicore interference can hurt the effective performance and energy consumption of the CPU (we discuss metrics to quantify this impact in Section 6). As a result, several approaches have been proposed to mitigate interference in the memory hierarchy (see Section 4.3).
@@
Modern LLCs typically make use of hierarchical designs to provide the large capacities required by contemporary workloads. A modern LLC can contain hundreds of thousands of cache sets, and if we were to maintain a monolithic array for the entire LLC capacity, we would have to deal with very large wire delays, as a bitline would potentially be shared by hundreds of thousands of rows, increasing the capacitance and resistance of both the wire itself and the parasitic impacts of each attached row. The increased capacitance and resistance translate into linear increases in delay and energy, and if left unaddressed can undermine the utility and efficiency of the LLC. Compounding the issue is the need for more cache ports, as multiple cores may want to access the cache at the same time, and a single-ported LLC can lead to starvation during periods of heavy cache request queuing.
To work around capacity and port limitations, modern caches rely on the concept of cache slicing, where a cache level is decomposed into multiple smaller pieces. [^10] Cache slicing has been referred to by several terms over the years, including cache interleaving and cache multi-banking. Early works on cache interleaving date to 1980 [76], and initial implementations interleaved (i.e., distributed) the words within a single cache block across multiple cache banks. In this work, we use the term cache slicing to refer to cache block interleaving across cache slices/banks, where a single slice/bank contains the entire cache block [80]. Each cache slice contains a subset of the total number of cache sets in the level, and is further partitioned by the number of ways. The slicing requires each cache access to first go through a slice decoder to identify which slice to look up. While early slice decoder implementations used a subset of the cache block address bits (e.g., using two of the cache index bits to decide which of four slices the block is assigned to), several commercial CPUs employ complex hash-based set mapping functions (e.g., [51]) to avoid hotspots (i.e., the uneven distribution of cache accesses across slices).
@@
For a given mix of workloads, a multicore CPU with n cores can potentially issue cache misses to main memory at as much as n times the rate that a single-core CPU does (note that this assumes similar core architectures in both CPUs, and depends heavily on the specific workloads and on the impact of cache interference between cores). This requires the main memory to scale up its ability to respond to these cache misses, and we see two examples of this in modern memory subsystems. First, more MSHRs are provided at the LLC in order to sustain a greater number of concurrent memory accesses. An LLC miss must stall any time the cache does not have a free MSHR available to allocate. If there are not enough MSHRs in the system, then such stalls are more likely to occur. As each cache slice often has its own set of MSHRs, multicore CPUs take advantage of the increased slice count to provide more MSHRs at a low cost, compared to scaling up a monolithic CAM. Second, the memory bandwidth has been scaling up to accommodate the increased demand. The memory bandwidth provided by DRAM has been increasing significantly over the last two decades, through a combination of (1) increasing bus frequencies for the memory channel and (2) the emergence of new DRAM types such as DDR5 [42] and High-Bandwidth Memory (HBM) [41]. While memory bandwidth can also be scaled by increasing the number of memory channels per CPU, multicore CPUs have largely avoided this approach due to the limited number of pins available in the CPU chip package.
The sharing of a single main memory across multiple cores can introduce new changes to the behavior of three properties, compared to their behavior with single-core CPUs. First, the optimal choice of an address interleaving scheme (i.e., which bits index which memory structures) can change depending on the type of workload executing on the CPU. Second, the optimal choice of row policy (i.e., whether a DRAM row is left open after all currently-queued requests are serviced) can depend on the type of workload as well. Third, the memory scheduling algorithm (i.e., the order in which queued DRAM requests are serviced) may need to change to avoid starving requests in certain scenarios. We discuss each of these in detail in Section 4.2.
@@ 3.3 Coordinating Memory Requests Across Cores
In order to maintain program correctness, multicore CPU must ensure that data updates are coordinated across all of the cores (as is the case in any parallel computer architecture). For the shared-memory programming model typically used for multicore CPUs, this involves two types of coordination [65]: (1) cache coherence, which ensures that updates to a single unit of data (e.g., one cache block) are made visible to and ordered across all cores; and (2) memory consistency, which ensures that updates across multiple units of data are interleaved according to a predetermined policy across all cores. We discuss cache coherence and memory consistency in more detail in Sections 4.4 and 4.5, respectively.
@@ 4 Managing Memory
The cores in a multicore CPU share memory with each other, as discussed in Section 3.2. This sharing introduces a number of issues that must be considered by both programmers and architects. We will discuss several of these issues in this section. To start, we define what it means to share memory, both from a software perspective and a hardware perspective. Then, we discuss main memory management, and how management policies in hardware can impact the performance of threads running on a multicore CPU. Finally, we discuss examples of cache coherence and memory consistency protocols that ensure correct program behavior in a shared-memory environment.
@@ 4.1 Shared-Memory Model
First, for a core, we need to define when a store is considered to be performed (i.e., the moment at which the stored value becomes visible to other cores). For example, for an out-of-order core, we can say that a store is performed once the store instruction is committed. Prior to commit, there are several events that can cause an executed store to be squashed (e.g., exception, misspeculation), making the commit event the earliest time we can guarantee that the store can safely proceed. As a result, we buffer stores inside the core, and do not initiate the store operation to the first-level cache until after the store instruction commits. A physical store buffer (sometimes called a write buffer) contains all committed writes that are yet to be issued to the cache. Depending on the memory consistency model, this store buffer may or may not be considered a part of shared memory; if it is, the store values in the buffer must be made available to other cores in a way that does not violate the consistency model (see Section 4.5 for more).
Second, the CPU must ensure that microarchitectural optimizations for memory operations obey a consistent set of rules, which are part of the instruction set architecture (ISA), so that programmers do not encounter program correctness issues during execution. All private and shared caches are considered part of the shared memory, which means that we need cache coherence to propagate the value of a cache store to other caches and cores, as we discuss in Section 4.4. Both within a thread and across threads, the interleaving of a store with other stores and with loads must obey a set of rules that we promise the programmer, which are defined as part of the memory consistency model. As we discuss in Section 4.5, some memory consistency models have strict interleaving expectations, which can simplify programming complexity at the expense of high performance overheads, while more relaxed memory consistency models can allow different cores to observe different interleavings (following a defined set of guarantees) in order to improve performance.
@@ 4.2 Main Memory Policies
Address interleaving (also known as memory interleaving) schemes [5] determine the way in which address bits are assigned to index different levels of the main memory organization, in an attempt to take advantage of spatial locality and memory-level parallelism (MLP). As shown in Figure 9, a typical DRAM-based main memory includes one or more memory channels, where each channel contains one or more DRAM modules. Within each module, there are multiple chips, with the chips grouped into one or more ranks. Each rank (a group of DRAM chips that operate in lockstep) contains multiple banks of DRAM, where each bank can mostly operate in parallel with each other. Inside each bank, at most one row of DRAM can be open at a time, where a row holds multiple columns of data (with each row in commonplace DRAM types containing a few kilobytes of data).
@@
Figure 11: Scheduling of multiple read requests when a bank conflict occurs (top), and when there is bank-level parallelism that avoids a bank conflict (bottom). We use the notation $B_{b}R_{r}C_{c}$ to indicate a memory access to Bank $b$, Row $r$, Column $c$. In this example, we assume that all requests access the same memory channel. The memory controller for the channel sends a request to the memory via the address, bus just before it is safe to initiate the operation inside memory (e.g., opening a row). The controller must ensure that operations to the same bank do not overlap and data transfers from memory to the controller on the shared data bus do not overlap.
Ultimately, a scheduling algorithm enforces these many constraints by prioritizing currently-queued requests based on a predefined policy and on the current state of the DRAM, and by enforcing a series of timing parameters that are declared as part of the specification of each DRAM type (e.g., DDR5 [42], HBM [41]). A commonly-implemented algorithm is known as first-ready, first-come first-serve (FR-FCFS) [70], which prioritizes requests to already-activated rows over older requests. FR-FCFS does this to improve average memory access time (AMAT), as the time required to activate a row, and to precharge (i.e., close) the row after requests are finished, can both take as long as the read and write requests themselves [42]. Memory controllers in commercial multicore CPUs continue to use FR-FCFS without introducing any notion of thread or application awareness [5], which can introduce interference across threads (see Section 4.3).

### v2

@@ 1 Introduction
This chapter will examine six topics related to multicore CPU architectures. First, it will motivate the benefits and limitations of processing multiple tasks concurrently, and how they drove the need for parallel processing. Second, it will examine the hardware design of a typical multicore CPU, and the changes required to support the efficient execution of multiple programs on multiple CPU cores. Third, it will study how memory management policies change to handle the increased traffic from multiple cores. Fourth, it will address software issues that optimize the use of multicore CPUs. Fifth, it will introduce common metrics that are used in the context of multicore CPUs. Finally, it will close by briefly discussing how commercial multicore CPUs have evolved since their introduction.
@@ 2 Motivating the Need for Concurrent Processing
As early as the days of analog computing, the benefits of parallelizing tasks became a clear goal: Luigi Federico Menabrea, in his 1842 study of Charles Babbage’s Analytical Engine [60], noted that a key advantage of mechanized computation would be its ability to produce several results at the same time. In the mid 20th century, as digital computers evolved rapidly, there was a pressing need to maximize both the performance and utilization of these computers, and the extraction of parallelism, in its various forms, became a key approach to meet these needs. While this chapter focuses on multicore CPUs, a brief taxonomy of parallel computing hardware is provided in Section 2.1.
@@ 2.1 Classifying Parallel Computing Hardware
Under Flynn’s taxonomy, non-pipelined, non-superscalar single-core CPUs that employ a von Neumann architecture are a canonical example of a SISD architecture. Examples of SIMD architectures include graphics processing units (GPUs), array processors, and vector extensions for CPUs. One reasonably-accepted example of a MISD architecture is redundant computing, such as the systems employed in avionics units. [^3] There is some latent controversy about which category pipelined CPUs fall into. In its best imitation of Switzerland and its neutrality, this chapter does not take a stand for any category. Sorry for being so bland. The MIMD category includes most computers that are capable of multiprocessing (Section 2.2), including computers with multicore CPUs.
In a multicore CPU, each CPU core is capable of executing one or more independent streams of instructions, with each instruction stream operating on its own data set (though this data set can potentially overlap with the data sets being used by other cores). To support this MIMD model of execution, each core has its own control logic, and much of the computation takes place independent from other cores. Section 3 elaborates on this more, while also discussing when coordination between cores does occur (both explicitly and implicitly).
@@ 2.3 Thread-Level Parallelism (TLP) Within an Application
where $f$ is the fraction of the program that can be parallelized, and $n$ is the number of concurrent threads that can execute the parallelizable part of the program. To visualize Amdahl’s Law, Figure 3 shows an example where $f=0.4$ (i.e., 40% of the program can be parallelized), for various thread counts. As the figure shows, for every doubling of the thread count, the reduction in total time is only half of the reduction achieved by the previous doubling.
@@
Putting this all together, Figure 5 shows a synthetically-constructed example of the parallel speedup one could expect to observe on a real parallel system. The example assumes that 99% of the application can be parallelized across all threads, while the remaining 1% must execute sequentially. It also assumes that the machine has 32 CPU cores, and since it has a fixed set of available resources, Amdahl’s Law can be used to predict performance. There are three observations from the figure. First, even with a single thread, the observed speedup is less than 1. This is because parallel speedup is compared against the best sequential implementation (see Section 6), and a parallel implementation (with a configurable parameter $n$ that sets the thread count) typically has overheads associated with adding code to support parallel execution, and these overheads are observed when $n=1$. Second, super-linear speedups can even exceed perfect parallelism. As mentioned above, this is the result of threads helping each other through data locality. Third, after a certain thread count (24 in our example), the performance at a higher thread count is actually lower than the performance at a lower thread count. This is because the benefits of additional thread-level parallelism and additional data locality are overcome by synchronization and interference overheads.
@@ 2.4 What to Do With All These Transistors?
Through the 1990s (and drawing some very broad, and likely inaccurate, generalizations), the vast majority of parallel computing architectures catered to the large-scale computing market, with much of the effort focused on supercomputers. Most personal computers (PCs) and workstations [^6] At that time, there was a distinction between more mainstream PCs and high-performance workstations that targeted power users. Today, in large part due to advancements in PC capabilities throughout the 1990s, this distinction has mostly disappeared, but the chapter references it here for historical context. incorporated only a single-core CPU (or in the case of some high-end workstations, multiple single-core CPUs installed on a multi-socket motherboard). Much of the focus on architectural innovation remained on improving the performance of the single core, leading to the maturation and widespread commercial adoption of techniques such as out-of-order processing and superscalar execution.
@@
As the breakdown of Dennard scaling materialized during the first few years of the 2000s, the power consumption issue was exacerbated significantly. This added pressure led the next wave of multicore CPUs to make a key trade-off to avoid the impending limits of thermal density: they operated each CPU core at a lower frequency than their single-core contemporaries, in the hopes that the increased parallel processing capability would make up for reduced single-core performance [20]. Section 3.1 discusses more about the CPU core trade-offs, which allowed architects to harness TLP instead of aggressive ILP to improve performance while keeping power issues under control, and opened the floodgates for the proliferation of multicore CPUs.
@@ 3.1 Optimizing CPU Cores for Parallelism
Multicore CPU designers prevent the design and verification of the CPU from increasing linearly with the core count by relying on design reuse through tiling. In a homogeneous multicore CPU, the design of the cores is identical, in order to reduce manufacturing and verification/testing costs. (Section 7 discusses popular approaches for heterogeneous multicore CPU design.) Figure 6 shows how the majority of the chip for a typical homogeneous multicore CPU is made up of multiple tiles, where each tile has exactly the same components. A tile usually consists of a single CPU core and the private caches associated with that core. Some architectures include a slice of the last-level cache in the tile (see Section 3.2). In addition to these tiles, the chip includes shared components (e.g., controllers that connect the CPU cores to external components, interconnects).
@@ 3.2 Sharing Caches and Main Memory
One issue with sharing resources across CPU cores is interference [6], where due to the finite amount of available resources, multiple cores may contend with each other for their desired share of these resources. Let us look at a contrived example, with a two-core CPU, where one core is servicing a cache miss that must retrieve data from main memory. As the data for that miss arrives at the LLC, the LLC must evict an existing cache block (assuming that no invalid blocks remain). Interference can occur when the eviction is for a cache block belonging to the other core, as that core will now miss on a subsequent access to the evicted block. Had the first core not triggered the eviction, this subsequent access would have been a cache hit, avoiding the long-latency miss to main memory. The resource contention that results from multicore interference can hurt the effective performance and energy consumption of the CPU (Section 6 discusses metrics to quantify this impact). As a result, several approaches have been proposed to mitigate interference in the memory hierarchy (see Section 4.3).
@@
Modern LLCs typically make use of hierarchical designs to provide the large capacities required by contemporary workloads. A modern LLC can contain hundreds of thousands of cache sets, and if a CPU designer were to maintain a monolithic array for the entire LLC capacity, they would have to deal with very large wire delays, as a bitline would potentially be shared by hundreds of thousands of rows, increasing the capacitance and resistance of both the wire itself and the parasitic impacts of each attached row. The increased capacitance and resistance translate into linear increases in delay and energy, and if left unaddressed can undermine the utility and efficiency of the LLC. Compounding the issue is the need for more cache ports, as multiple cores may want to access the cache at the same time, and a single-ported LLC can lead to starvation during periods of heavy cache request queuing.
To work around capacity and port limitations, modern caches rely on the concept of cache slicing, where a cache level is decomposed into multiple smaller pieces. [^10] Cache slicing has been referred to by several terms over the years, including cache interleaving and cache multi-banking. Early works on cache interleaving date to 1980 [77], and initial implementations interleaved (i.e., distributed) the words within a single cache block across multiple cache banks. This chapter uses the term cache slicing to refer to cache block interleaving across cache slices/banks, where a single slice/bank contains the entire cache block [81]. Each cache slice contains a subset of the total number of cache sets in the level, and is further partitioned by the number of ways. The slicing requires each cache access to first go through a slice decoder to identify which slice to look up. While early slice decoder implementations used a subset of the cache block address bits (e.g., using two of the cache index bits to decide which of four slices the block is assigned to), several commercial CPUs employ complex hash-based set mapping functions (e.g., [52]) to avoid hotspots (i.e., the uneven distribution of cache accesses across slices).
@@
For a given mix of workloads, a multicore CPU with n cores can potentially issue cache misses to main memory at as much as n times the rate that a single-core CPU does (note that this assumes similar core architectures in both CPUs, and depends heavily on the specific workloads and on the impact of cache interference between cores). This requires the main memory to scale up its ability to respond to these cache misses, and modern memory subsystems have two examples of this. First, more MSHRs are provided at the LLC in order to sustain a greater number of concurrent memory accesses. An LLC miss must stall any time the cache does not have a free MSHR available to allocate. If there are not enough MSHRs in the system, then such stalls are more likely to occur. As each cache slice often has its own set of MSHRs, multicore CPUs take advantage of the increased slice count to provide more MSHRs at a low cost, compared to scaling up a monolithic CAM. Second, the memory bandwidth has been scaling up to accommodate the increased demand. The memory bandwidth provided by DRAM has been increasing significantly over the last two decades, through a combination of (1) increasing bus frequencies for the memory channel and (2) the emergence of new DRAM types such as DDR5 [43] and High-Bandwidth Memory (HBM) [42]. While memory bandwidth can also be scaled by increasing the number of memory channels per CPU, multicore CPUs have largely avoided this approach due to the limited number of pins available in the CPU chip package.
The sharing of a single main memory across multiple cores can introduce new changes to the behavior of three properties, compared to their behavior with single-core CPUs. First, the optimal choice of an address interleaving scheme (i.e., which bits index which memory structures) can change depending on the type of workload executing on the CPU. Second, the optimal choice of row policy (i.e., whether a DRAM row is left open after all currently-queued requests are serviced) can depend on the type of workload as well. Third, the memory scheduling algorithm (i.e., the order in which queued DRAM requests are serviced) may need to change to avoid starving requests in certain scenarios. Section 4.2 discusses each of these in detail.
@@ 3.3 Coordinating Memory Requests Across Cores
In order to maintain program correctness, multicore CPU must ensure that data updates are coordinated across all of the cores (as is the case in any parallel computer architecture). For the shared-memory programming model typically used for multicore CPUs, this involves two types of coordination [66]: (1) cache coherence, which ensures that updates to a single unit of data (e.g., one cache block) are made visible to and ordered across all cores; and (2) memory consistency, which ensures that updates across multiple units of data are interleaved according to a predetermined policy across all cores. Sections 4.4 and 4.5 discuss cache coherence and memory consistency, respectively, in more detail.
@@ 4 Managing Memory
The cores in a multicore CPU share memory with each other, as discussed in Section 3.2. This sharing introduces a number of issues that must be considered by both programmers and architects. This section will discuss several of these issues. To start, it defines what it means to share memory, both from a software perspective and a hardware perspective. Then, it discusses main memory management, and how management policies in hardware can impact the performance of threads running on a multicore CPU. Finally, it discusses examples of cache coherence and memory consistency protocols that ensure correct program behavior in a shared-memory environment.
@@ 4.1 Shared-Memory Model
First, for a core, one needs to define when a store is considered to be performed (i.e., the moment at which the stored value becomes visible to other cores). For example, for an out-of-order core, a popular definition is that a store is performed once the store instruction is committed. Prior to commit, there are several events that can cause an executed store to be squashed (e.g., exception, misspeculation), making the commit event the earliest time that the store is guaranteed to safely proceed. As a result, stores are buffered inside the core, and the store operation to the first-level cache is not initiated until after the store instruction commits. A physical store buffer (sometimes called a write buffer) contains all committed writes that are yet to be issued to the cache. Depending on the memory consistency model, this store buffer may or may not be considered a part of shared memory; if it is, the store values in the buffer must be made available to other cores in a way that does not violate the consistency model (see Section 4.5 for more).
Second, the CPU must ensure that microarchitectural optimizations for memory operations obey a consistent set of rules, which are part of the instruction set architecture (ISA), so that programmers do not encounter program correctness issues during execution. All private and shared caches are considered part of the shared memory, which means that cache coherence is needed to propagate the value of a cache store to other caches and cores, as Section 4.4 discusses. Both within a thread and across threads, the interleaving of a store with other stores and with loads must obey a set of rules that are promised to the programmer, which are defined as part of the memory consistency model. As Section 4.5 discusses, some memory consistency models have strict interleaving expectations, which can simplify programming complexity at the expense of high performance overheads, while more relaxed memory consistency models can allow different cores to observe different interleavings (following a defined set of guarantees) in order to improve performance.
@@ 4.2 Main Memory Policies
Address interleaving (also known as memory interleaving) schemes [5] determine the way in which address bits are assigned to index different levels of the main memory organization, in an attempt to take advantage of spatial locality and memory-level parallelism (MLP) [28]. As shown in Figure 9, a typical DRAM-based main memory includes one or more memory channels, where each channel contains one or more DRAM modules. Within each module, there are multiple chips, with the chips grouped into one or more ranks. Each rank (a group of DRAM chips that operate in lockstep) contains multiple banks of DRAM, where each bank can mostly operate in parallel with each other. Inside each bank, at most one row of DRAM can be open at a time, where a row holds multiple columns of data (with each row in commonplace DRAM types containing a few kilobytes of data).
@@
Figure 11: Scheduling of multiple read requests when a bank conflict occurs (top), and when there is bank-level parallelism that avoids a bank conflict (bottom). The notation $B_{b}R_{r}C_{c}$ indicates a memory access to Bank $b$, Row $r$, Column $c$. This example assumes that all requests access the same memory channel. The memory controller for the channel sends a request to the memory via the address, bus just before it is safe to initiate the operation inside memory (e.g., opening a row). The controller must ensure that operations to the same bank do not overlap and data transfers from memory to the controller on the shared data bus do not overlap.
Ultimately, a scheduling algorithm enforces these many constraints by prioritizing currently-queued requests based on a predefined policy and on the current state of the DRAM, and by enforcing a series of timing parameters that are declared as part of the specification of each DRAM type (e.g., DDR5 [43], HBM [42]). A commonly-implemented algorithm is known as first-ready, first-come first-serve (FR-FCFS) [71], which prioritizes requests to already-activated rows over older requests. FR-FCFS does this to improve average memory access time (AMAT), as the time required to activate a row, and to precharge (i.e., close) the row after requests are finished, can both take as long as the read and write requests themselves [43]. Memory controllers in commercial multicore CPUs continue to use FR-FCFS without introducing any notion of thread or application awareness [5], which can exacerbate interference across threads (see Section 4.3).